OpenStudy (anonymous):

1/2y + 1/4x = 2.5 in slope-intercept form? (Y= mx+ b)

OpenStudy (blurbendy):

1/2 y = -1/4x + 2.5 y = -1/4*2 x + 2.5*4 Can you solve from there?

OpenStudy (blurbendy):

oops 2.5*2

OpenStudy (anonymous):

why would it be Negative 1/4 x

OpenStudy (blurbendy):

It's negative 1/4 because we subtracted 1/4x from both sides in order to get the equation in the form y = mx + b. -1/4 is our new "m"

OpenStudy (anonymous):

Okay one second. Let me solve this.

OpenStudy (blurbendy):

Alright.

OpenStudy (anonymous):

y = -1/4 x + 5 is the simplfyed equation..?

OpenStudy (anonymous):

*simplified

OpenStudy (blurbendy):

Close, -1/4 * 2 = -1/4 * 2 / 1

OpenStudy (anonymous):

y = -2/4 x + 5

OpenStudy (anonymous):

..?

OpenStudy (blurbendy):

you're so close, -2/4 can be simplified a little bit more. If you have 4 slices of pie, and you eat two slices, you've eaten _____ of the pie.

OpenStudy (anonymous):

xD Oh god. I'm sorry. y = -1/2 x + 5

OpenStudy (blurbendy):

Nice, you got it!

OpenStudy (anonymous):

Okay and what about y + 1/4x = 2?

OpenStudy (anonymous):

Is it Y=-1.4x + 2

OpenStudy (blurbendy):

yeah! just subtract -1 /4x like last time to get it into the form y = mx + b.

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (blurbendy):

No problem! Good luck :)